OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)? Answer y = – 2x + 9 y = 2x – 11 y = – 2x + 6 y = 2x – 8

OpenStudy (katrinakaif):

First we have to find the slope: y2-y1 ---- x2- x1 Your given points are: (2,-7) or (x1,y1) (5,-1) or (x2,x1) Once you have found "m" or slope, choose any one of the given points and then plug it in into the formula: y-y1 = m(x-x1) and simplify

OpenStudy (anonymous):

thnks

OpenStudy (katrinakaif):

Anytime. =)

OpenStudy (anonymous):

i got the third choice is it right?

OpenStudy (katrinakaif):

One moment.

OpenStudy (katrinakaif):

- 1 + 7 ------ 5 - 2 6 / 3 = 2 Your slope is incorrect, so no. Try again

OpenStudy (anonymous):

oh ok its the last one!

OpenStudy (katrinakaif):

Which set of points did you use?

OpenStudy (anonymous):

wait im confused!

OpenStudy (katrinakaif):

Which set of points did you use to plug into the equation: y-y1 = m (x-x1)